A body A starting from rest moves with an acceleration a on a straight line. Body B starts after some time t from the same point and moves on the same path with a constant velocity v. If B meets A just once on its path, what is the velocity of A when they meet?

Grade:10

2 Answers

KAPIL MANDAL
6 years ago
Let`s say they meet at time t. Distance traversed by A in time T is S=(1/2)*a*T^2.Since they meet, distance traversed by B is also same. Also,for B, S= v (T-t). Hence, (1/2)*a*T^2 = v (T-t). Solve T from this equation and put it in the formula, V=aT
KAPIL MANDAL
6 years ago
Sorry, the 1st line is Let`s say they meet at time T. Everything else is same. Plz ask if you hv any doubt